# Designing Guiding Principles for NLP for Healthcare: A Case Study of Maternal Health

### 19 Dec 2023 v1

@@ 4 RESULTS
19% HC management/admin
16% HC services researcher
11% other peri. HC provider
11% other non-HC peri. support
@@
Table 1: Demographic description of participant cohorts. Healthcare workers and birthing people were recruited from the Prolific platform while the workshop participants took part in multiple training and educational sessions to learn about LLMs, their applications to maternal health, and their risks. Work category, race/ethnicity, highest education, and gender all allowed multiple selections, so the percents for these categories might not sum to 100%.
4.1 Survey Results
We release our survey results publicly to support future research. [^3] We remove demographic features for the workshop participants, as these participants interacted with one another and could potentially identify others’ survey responses. Participants consented to the sharing of their responses. All survey data is available in supplementary materials and at this public Github repository: https://github.com/maria-antoniak/maternal-health-principles Table 1 provides a summary of participant demographics and background. The workshop participants tended to represent the non-profit, community health, research, and governmental sectors in contrast to clinicians such as physicians and nurses that were more common in the Prolific cohort of healthcare workers. The workshop participants represented a more even distribution across racial/ethnic groups, but all cohorts had little representation in the East Asian, South Asian, and Southeast Asian groups. Overall, workshop participants reported less experience with AI and NLP than birthing people and healthcare workers recruited from Prolific (see Appendix A).
Figure 4 shows how frequently each value was selected by each cohort. Overall, safety, privacy, and performance were selected more often by birthing people and healthcare workers in comparison to workshop participants, who were more likely to select inclusiveness. The birthing people and workshop participants were united in being more likely to select human autonomy than the healthcare workers, and less likely than the healthcare workers to select transparency and accountability. Figure 5 compares coocurrences of values selected by different cohorts. Birthing people were more consistent in their selections, while healthcare workers had more variation in their selections. Overall, we do not find large differences in the value selections across stratifications for generalized trust, trust in healthcare workers, or generalized AI perceptions (see Appendix A).
Figure 4: Frequencies of selected values. Each participant was asked to select five values from a list of 12 curated by Jakesch et al. [20]. Participants were given definitions of the values, also drawn from Jakesch et al. [20]. Birthing people and healthcare workers overall responded more similarly to each other than to the workshop participants, but this was not always the case (e.g., transparency, human autonomy).
Figure 5: Co-occurrences of selected values. Each participant was asked to select five values from a list of 12 curated by Jakesch et al. [20]. Participants were given definitions of the values, also drawn from Jakesch et al. [20]. Overall, the workshop participants were less consistent in their selection of values than the birthing people and healthcare workers.
4.2 Thematic analysis of discussion sessions
@@
We identified themes and concepts frequently brought up by study participants in workshop breakout discussions, written comments, free-text survey answers, and chatbot queries. We first performed open coding on responses from each source to identify relevant concepts, then axial coding to group these concepts under broad themes [10]. The authors collectively discussed these themes and iterated several times before synthesizing them into the final set of nine guiding principles presented in §4.3. We summarize these principles in the following section and include anonymized quotes from our study participants to highlight stakeholder pain points and further support the inclusion of each principle. Additionally, we provide actionable recommendations to NLP researchers and practitioners based on our own experience and expertise as NLP researchers and healthcare workers (Appendix B). Our thematic analyses also raised several additional points, which we discuss in §5.
4.3 Guiding Principles
@@
Drawing from the discussion themes, our survey results, and related work (both applications work and high-level ethics guidelines like those by Chen et al. [9] and D’Ignazio and Klein [15]), we develop guiding principles for the responsible use of NLP in maternal health. Detailed descriptions and actionable recommendations derived from each principle are provided in Appendix B.
@@
We present these principles grouped under three themes: context, measurements, and values. Context principles ask stakeholders to incorporate the fundamentals and history of maternal health in their applications, measurement principles discuss what to optimize for and how to evaluate, and value principles address how NLP researchers and practitioners should situate user voices and data relevant to our systems. We summarize each principle below and highlight in participants’ own words their importance. In quotes, workshop participants are attributed as W1-10, birthing people as B1-11, and healthcare workers as H1-15.
Theme 1: Context
@@
Be aware of power dynamics in the care team  Maternal healthcare has a long and fraught history of shifting power dynamics in the care team [4]. These historical shifts, like male OB/GYNs replacing midwives in the early 20th century, can lead to both good and bad effects for the birthing person. NLP practitioners should know this history and work to improve rather than exacerbate existing hierarchies, being mindful of the impact of tools and automation on midwives, doulas, and other care workers whose placement is already precarious. “AI should not be used instead of actual and necessary medical care” (B1). “If mothers relied too heavily on AI instead of seeking professional help then the nurses and doulas may see fewer people seeking care” (B11).
@@
Know the politics and implications of your measurements  Using NLP tools for data collection can have unintended consequences, e.g., (i) overemphasis on data collection can draw resources away from building solutions, (ii) reproductive health is politicized, and measurements may be used as evidence in support of unexpected agendas, and (iii) focusing only on measuring problems can contribute to “deficit narratives” that blame communities for their own challenges. “It’s key that the AI is actually able to provide solutions to problems that can be fixed, and not just simply acknowledging them” (H4). We therefore recommend clarifying the impact goal of an NLP project before focusing on what data to collect. Financial motivations also muddy these decisions. As one participant puts it, “I get worried about what’ll happen when insurance companies think there’s cost savings to using these tools…they can cut corners, have more profit…given the incentives in … the healthcare system” (W1).
@@
Learn from healthcare traditions and communities  Thousands of years of cultural traditions, healthcare practices, and tool development already exist, and more recently, grassroots communities on the internet have formed to help birthing people prepare for and work through their challenges and experiences. “People have the mindset that this needs to actually replace things or that this needs to be the way in the future of addressing maternal health…diminishing all the work that’s been done in the past and the positives of having people that truly understand this work” (W2). Learning from traditions and communities can help avoid repeating mistakes made by others. “It [(AI)] should follow the same guidelines that medical professionals do: ‘Do no harm’ ” (H6).
Theme 2: Measurements
@@
Optimize for outcomes that support the whole person  “It’s not just about the outcome, right? It’s about the whole experience” (W3). NLP research, like most AI research, is centered on structured tasks with defined and accessible outcomes, like maternal mortality rates. Reducing maternal mortality is important, but so are other outcomes, like achieving a positive birthing experience or breastfeeding without stress. As one participant says, “it would be great if AI could tell everything from conception to birth” (B10). Taking a more expansive view [^4] By “expansive view,” we do not mean that a single model must be capable of all tasks (as this may in fact be a weakness of current LLMs whose generalizability can lead to a lack of clear evaluation). Rather, we invite NLP researchers and practitioners to be open-minded about tasks they can address with their NLP work. of the tasks and optimizations that could benefit maternal health stakeholders can open the door to other opportunities. “A dream AI tool should have the ability to address and develop solutions for mothers who are struggling to cope with the demands of pregnancy and overall maternal health” (H4).
@@
Fight against inequity and protect all groups  The healthcare system treats people differently on the basis of race, class, gender, etc. NLP models can mirror biases found in training datasets—and even worse, they can over-represent and exacerbate the impacts of those biases. “Everything that has the potential to benefit also has the potential to hurt” (W4). When evaluating NLP systems, one should consider both disparities in impact (health outcomes for different groups) and disparities in treatment (whether people with the same complaint receive different treatment). Tools should “allow providers to improve their own cultural competency or race-consciousness” (W5), and empower “those lacking social and financial support,” such as young mothers (H14).
@@
Hold onto the human: empathy, emotion, relationships, complexity  NLP tools should account for human elements like emotion, empathy, and storytelling [8], remembering that each birth is unique and each birthing person has a unique set of circumstances, experiences, and preferences. “Human connection is important and should be emphasized so that better quality service can be provided” (W6); “your own judgment and/or human compassion components, wisdom, experience [are a] part of the care” (W7). Don’t shy away from outliers and difficult, subjective tasks. Sometimes the goal is not to solve a problem, but to better understand it. “Birth is so complicated though that I don’t see women ever replacing medical care with AI advice” (B11).
Theme 3: Values
@@
Include the voices of those seeking care  The voices of the most vulnerable stakeholder group, birthing people, should be included in the design and development of NLP tools. While it may not always be possible to include birthing people on a research or production team, it is important to integrate their perspective throughout design and deployment processes through surveys and user studies. As one participant puts it, “there needs to be community input, there needs to be representation in creating these tools” (W8).
@@
Always center the agency and autonomy of the person seeking care  Maternal healthcare has an unfortunate history of abuse and disregard for the birthing person’s agency. Rather than working from a perspective of skepticism about the birthing person’s ability to make decisions for themselves, practitioners should respect the person’s agency over their own experience. How NLP tools are being used to make decisions in maternal healthcare should be disclosed to the birthing person, such that “the person using it knows what kind of information or advice it can and cannot give” (B2). Correspondingly, some tools should not be built, such as anything “infringing on the rights of the mom or baby” (B3).
@@
Respect and support your data sources  Generations of birthing people have passed down oral stories, written books, and created online content about pregnancy, labor, and the postpartum period. NLP studies and tools benefit from this knowledge and data, and should give credit and be designed to avoid supplanting systems of support that are already thriving. As several workshop participants observed, ChatGPT and other similar chatbots “don’t really note their references of where they’re getting the information from” (W9). “The principle that should guide these tools is to… have transparency for its sourcing of data” (H2).
5 DISCUSSION
We asked over 100 people from diverse groups about their values, practices, and perceptions of AI and chatbots. Using maternal health as a test case, we uncovered a set of guiding principles with lessons for all uses of NLP for healthcare. Our discovered themes touch on questions of power and agency, trust between health specialists and birthing people, and contextualizing innovations within a history of disenfranchisement and mistakes. Our study also opened further questions, which we discuss below.
Applicability to other healthcare settings
Maternal health served us as a representative health case study through which we could apply a focused lens to the use of LLMs for healthcare. Related healthcare settings include resource shortages in primary care [7], individuals seeking health information to diagnose symptoms [2], and racial considerations in disease severity models [29]. In each case, our guiding principles can help stakeholders think through whether and how to incorporate NLP tools into healthcare settings.
Weighing the risks and benefits of using NLP and LLMs in healthcare settings
Participants reported largely positive attitudes towards NLP tools with important caveats. Workshop participants unanimously agreed that NLP tools for maternal health should always be designed to be assistive rather than autonomous (holding on to the human) while having high hopes about the development and adoption of NLP tools in healthcare.
@@ 5 DISCUSSION
Settings in which NLP tools were viewed to be particularly useful to healthcare providers included: (i) reducing administrative burden, including better/faster communication and coordination of care across medical departments, and (ii) improving medical education and training, including assisting staff in developing better cultural sensitivity. On the other hand, tools for decision-making settings were viewed negatively. These results contrast with current trends in NLP research (§2), which tend towards surveillance of birthing people and tasks like risk prediction.
@@
Birthing people viewed NLP tools very positively, especially for providing information and recommendations (centering their agency and autonomy). For example, participants wished they had the chatbot during their pregnancy: “I wish it [(AI)] had been around when my son was a newborn so I could interact with it during late night feedings. One, to give me something to do, and two, to make me feel like I wasn’t alone” (B12). In particular, birthing people emphasized how comforting it would be to have a fast and convenient resource to assuage fears about whether what they are experiencing is normal: “Just to have something there to ask questions to when I am not sure as to what is happening or when I need a quick answer” (B9). However, some participants cautioned that such tools should provide additional context to help patients understand whether the suggestions or advice are applicable to their specific situation due to differing levels of medical understanding.
Improving the utility of current LLMs
While study participants were impressed the chatbot’s responses and capabilities, they identified several areas for improvement, in terms of both response presentation and response content. In addition to wanting more attribution of data sources (respecting data sources), participants thought response presentation could be improved by (i) reducing constant use of “safeguards,” which hampers readability (e.g., “I am an LLM…”, “please consult a doctor…”, etc.) and (ii) tailoring responses to match a user’s literacy and reading levels, particularly when presenting complex medical recommendations. This further motivates recent work on generating plain language summaries [1] and personalized rewrites of medical information [18], and points away from a general-purpose chatbot. Discussants also suggested (i) incorporating cultural nuance in model responses (e.g., personalizing responses for the user’s cultural context and away from a single dominant culture [32, 33]) and (ii) improving model sensitivity to minor changes in wording, which often generated very different responses.
LLMs as part of a larger ecosystem to meet information needs

### 23 Jan 2024 v2

@@ 3.1. Survey Results
19% healthcare management/admin
16% healthcare services researcher
11% other perinatal healthcare provider
11% other non-healthcare perinatal support
@@
Table 1. Demographic description of participant cohorts. Healthcare workers and birthing people were recruited from the Prolific platform while the workshop participants took part in multiple training and educational sessions to learn about LLMs, their applications to maternal health, and their risks. Work category, race/ethnicity, highest education, and gender all allowed multiple selections, so the percents for these categories might not sum to 100%.
3.2. Thematic Analysis of Discussion Sessions
@@
We identified themes and concepts frequently brought up by study participants in workshop breakout discussions, written comments, free-text survey answers, and chatbot queries. We first performed open coding on responses from each source to identify relevant concepts (Charmaz, 2014), then axial coding to group these concepts under broad themes (Corbin and Strauss, 1990). The authors collectively discussed these themes and iterated several times before synthesizing them into the final set of three themes and nine guiding principles presented in §4. We summarize these principles in the following section and include anonymized quotes from our study participants to highlight stakeholder pain points and further support the inclusion of each principle. Additionally, we provide actionable recommendations to NLP researchers and practitioners based on our own experience and expertise as NLP researchers and healthcare workers. Our thematic analyses also raised several additional points, which we discuss in §5.
4. Guiding Principles
@@
Drawing from the discussion themes, our survey results, and related work (both applied work and high-level ethics guidelines like those by Chen et al. (2021) and D’Ignazio and Klein (2023)), we develop guiding principles for the responsible use of NLP in maternal health.
@@
We present these principles grouped under three themes: context, measurements, and values. Context principles ask practitioners to incorporate the fundamentals and history of maternal health in their applications, measurement principles discuss what to optimize for and how to evaluate, and value principles address how practitioners should situate user voices and data relevant to their systems.
We summarize each principle below and highlight in participants’ own words their importance. In quotes, workshop participants are attributed as W1-10, birthing people as B1-11, and healthcare workers as H1-15.
4.1. Theme 1: Context
@@
Be aware of power dynamics in the care team.   D’Ignazio and Klein (2023) have argued that power should be a central concern of feminist data science, and maternal healthcare has a long and fraught history of shifting power dynamics in the care team (Benjamin, 2022). These historical shifts, e.g., male OB/GYNs replacing midwives in the early 20th century (Brodsky, 2008; Loudon, 2008), have led to the marginalization of some maternal health workers and both good and bad consequences for birthing people. NLP practitioners should know this history and work to improve rather than exacerbate existing hierarchies, being mindful of the impact of tools and automation on midwives, doulas, and other care workers whose placement is already precarious. For example, one of the healthcare workers in our study mentioned that “an AI chatbot can eliminate the need for a Doula” (H1), highlighting the vulnerabilities of certain workers. One of the birthing people framed this as a harm rather than an opportunity: “If mothers relied too heavily on AI instead of seeking professional help then the nurses and doulas may see fewer people seeking care” (B11).
Recommended actions for practitioners: Understand historical power shifts in the care team. Consider carefully who you expect to use your tool and who will be impacted (or potentially replaced) by your tool. Rather than designing a chatbot as a replacement for a core member of the care team, consider designing supportive tools that can improve the care team’s collaboration. Consider these questions: Whose language and data are represented in the training data of the tool? Who will receive predictions and advice? Who will have access to and control? Are you concentrating more power within a single care team role?
@@
Know the politics and implications of your measurements.   Using NLP tools for data collection can have unintended consequences that can impact the safety and agency of birthing people. For example, (i) overemphasis on data collection can draw resources away from building solutions (D’Ignazio and Klein, 2023), (ii) reproductive health is politicized, and measurements may be used as evidence in support of unexpected agendas, and (iii) focusing only on measuring problems can contribute to “deficit narratives” that blame communities for their own challenges (D’Ignazio and Klein, 2023). In our literature review (§1), we found that much prior NLP research for maternal health has focused on data collection and surveillance of birthing people, and participants pushed back against this pattern: “It’s key that the AI is actually able to provide solutions to problems that can be fixed, and not just simply acknowledging them” (H4). Financial motivations also muddy these decisions, with the healthcare industry often focused on reducing costs (Obermeyer and Mullainathan, 2019); as one participant put it, “I get worried about what’ll happen when insurance companies think there’s cost savings to using these tools…they can cut corners, have more profit…given the incentives in … the healthcare system” (W1).
Recommended actions for practitioners: Clarify your research and/or impact goals before collecting data. Consider the current allocation of resources and where help is most needed when choosing research questions. While additional data collection can confirm well-known problems, it can also draw resources away from other important problems. Think through possible narratives involved in public portrayals of your work, and invest time in framing your NLP and especially LLM outputs for better public understanding of risks (e.g., hallucinations) involved in these technologies. Know that your measurements and predictions could contain information that could be used to incriminate the birthing person in some places and situations. Similarly, your data and measurement methods could be used for unintended purposes, like personalized advertising and the setting of insurance rates.
@@
Learn from maternal health traditions and communities.   Thousands of years of cultural traditions, healthcare practices, and tool development related to maternal health already exist (Goode and Katz Rothman, 2017), and more recently, grassroots communities on the internet have formed to help birthing people prepare for and work through their challenges and experiences (Antoniak et al., 2019; Andalibi and Garcia, 2021). “People have the mindset that this needs to actually replace things or that this needs to be the way in the future of addressing maternal health…diminishing all the work that’s been done in the past and the positives of having people that truly understand this work.” (W2) Learning from traditions and communities can help avoid repeating mistakes made by others. “It [(AI)] should follow the same guidelines that medical professionals do: ‘Do no harm.’ ” (H6)
Recommended actions for practitioners: Combining NLP expertise with an interdisciplinary team can help you avoid reinventing the wheel. Consider how NLP tools can support and learn from storytelling, which is an important language-based way for communities and generations to collectively teach and learn about maternal health, as well as a way for birthing people to process their experiences (Kay et al., 2017; Carson et al., 2017). Consider how to sample training datasets without dismissing the care-seeker’s perspective; weigh the risks of possible misinformation against the value of personal experiences within your specific healthcare setting.
4.2. Theme 2: Measurements
@@
Optimize for outcomes that support the whole person. “It’s not just about the outcome, right? It’s about the whole experience” (W3). Most NLP research is centered on structured tasks with defined and accessible outcomes, e.g., predicting postpartum depression. Such tasks are important, but so are other outcomes, like achieving a positive birthing experience (Karlström et al., 2015) or breastfeeding without stress (Spannhake et al., 2021). As one participant says, “it would be great if AI could tell everything from conception to birth” (B10). Taking a more expansive view of the tasks and optimizations that could benefit maternal health stakeholders can open the door to other opportunities. By “expansive view,” we do not mean that a single model must be capable of all tasks (as this may in fact be a weakness of current LLMs whose purported generalizability can lead to a lack of clear evaluation (Raji et al., 2021)). Rather, we invite NLP researchers and practitioners to be open-minded about tasks they can address with their NLP work. “A dream AI tool should have the ability to address and develop solutions for mothers who are struggling to cope with the demands of pregnancy and overall maternal health” (H4).
Recommended actions for practitioners: Curate and create text datasets with novel annotations, valuing this data work (Sambasivan et al., 2021). Use social media and birthing person interviews to get a holistic overview of what outcomes birthing people care about (Andalibi and Garcia, 2021). Run long-term user studies to support a broader range of outcomes of interest. Consider including subjective and/or qualitative measurements.
@@
Protect all groups of birthing people.   The healthcare system treats people differently on the basis of race, class, gender, etc. (Smedley et al., 2003; Zavala et al., 2020; Hooper et al., 2020). NLP models can mirror biases found in training datasets (Caliskan et al., 2016; Steed et al., 2022), and even worse, they can over-represent and exacerbate the impacts of those biases (Zhao et al., 2017; Hall et al., 2022). “Everything that has the potential to benefit also has the potential to hurt” (W4). When evaluating NLP systems, one should consider both disparities in impact (health outcomes for different groups) and disparities in treatment (whether people with the same complaint receive different treatment). Tools should “allow providers to improve their own cultural competency or race-consciousness” (W5), and empower “those lacking social and financial support,” such as young birthing people (H14).
Recommended actions for practitioners: Recruit birthing people from diverse groups to participate and/or give feedback on your NLP study/tool. Consider the role of personalization in your NLP/LLM-based system, considering both disparate impacts and disparate treatments. Sometimes people with the same complaint are not given the same treatment (e.g., pain management disparities (Samulowitz et al., 2018)). But in other cases, applying the same treatment, without regard for individual circumstances, is also an inequity. When building LLM-based tools that communicate with birthing people (e.g., chatbots), evaluate across a diversity of vocabularies, accents, and languages (Mehandru et al., 2022). Beware counterfactuals as a methodological bandaid (Kasirzadeh and Smart, 2021). Measuring and addressing disparities are important, but simply switching tokens indicating race or gender (e.g., replacing all mentions of she with he) ignores society’s multi-dimensional impacts on people living with those identities, which can lead to correlated differences in text datasets.
@@
Hold onto the human: empathy, emotion, relationships, complexity.   Emotion, empathy, and storytelling are important components of healthcare (Charon, 2001), and NLP tools should account for such human elements, remembering that each birth is unique and each birthing person has a unique set of circumstances, experiences, and preferences. “Human connection is important and should be emphasized so that better quality service can be provided” (W6); “your own judgment and/or human compassion components, wisdom, experience [are a] part of the care” (W7). Working with language data, as opposed to structured datasets, opens doorways to subjective tasks such as measuring empowerment (Njoo et al., 2023), agency (Sap et al., 2017), and sentiment (Ali et al., 2013). “Birth is so complicated though that I don’t see women ever replacing medical care with AI advice” (B11).
Recommended actions for practitioners: Look for outliers and value what they can teach you, rather than removing them. A one-size-fits-all approach is probably not appropriate; prioritize personalization and/or allow the tool to be tailored by the individual birthing person or healthcare worker. Subjective language tasks like measuring framing (Card et al., 2015) can illuminate aspects of healthcare experiences overlooked by non-text-based methods. Include qualitative methods in your study design, complementing NLP models with interviews, grounded theory (Corbin and Strauss, 1990), and methods from the social sciences.
4.3. Theme 3: Values
@@
Include the voices of those seeking care.   The voices of the most vulnerable stakeholder group, birthing people, should be included in the design and development of NLP tools. Center the principle of “nothing about us without us,” most recently popularized by disability activists (Charlton, 1998). While it may not always be possible to include birthing people on a research or production team, it is important to integrate their perspective throughout design and deployment processes through surveys and user studies. As one participant put it, “there needs to be community input, there needs to be representation in creating these tools” (W8). At the same time, practitioners should be careful to engage fully with these voices, rather than using them as an ethical veneer on otherwise exploitative tools (Sloane et al., 2022).
Recommended actions for practitioners: Where possible, include birthing people and healthcare professionals in your research team. Include surveys and user studies throughout your research and design process. Incorporate literature written by birthing people, and learn from related work about the birthing experience and birthing people’s perspectives and needs. Be gender inclusive: do not automatically predict the gender of study subjects (Keyes, 2017; Larson, 2017), and remember to include the concerns of trans birthing people in your study design. Use gender inclusive language and follow the HCI Guidelines for Gender Equity and Inclusivity (Scheuerman et al., 2020).
@@
Always center the agency and autonomy of the birthing person.   Maternal healthcare has an unfortunate history of abuse and disregard for the birthing person’s agency (see §1.3). NLP practitioners, even when studying topics like misinformation, should not work from a perspective of skepticism about the birthing person’s ability to make decisions for themselves. How NLP tools are being used to make decisions in maternal healthcare should be disclosed to the birthing person, such that “the person using it knows what kind of information or advice it can and cannot give” (B2). Correspondingly, some tools should not be built (Bender et al., 2021), such as anything “infringing on the rights of the mom or baby” (B3).
Recommended actions for practitioners: Explore the construction of NLP tools that increase the agency available to the birthing person rather than making decisions for them. For example, NLP tools might be used to provide birthing people with additional decision points, resources, explanations, descriptions of what to expect, or assistance for communicating with healthcare providers. If NLP tools are used to make or assist in clinical decisions, this should be disclosed to birthing people, and where possible, providers should obtain direct consent from birthing people.
@@
Respect and support your data sources.   Generations of birthing people have passed down oral stories, written books, and created online content about pregnancy, labor, and the postpartum period. NLP studies and tools benefit from this knowledge and data and should give credit and be designed to avoid supplanting systems of support that are already thriving. As several workshop participants observed, ChatGPT and other similar chatbots “don’t really note their references of where they’re getting the information from” (W9). “The principle that should guide these tools is to… have transparency for its sourcing of data” (H2).
Recommended actions for practitioners: Give credit to the data sources. Use proper attribution, as this not only respects people’s work but also builds trust from your users and supports auditing of your system. Maintain privacy. Collect only necessary data and store the data securely. Avoid perpetuating the over-surveillance of birthing people. Beware the “paradox of reuse” (McMahon et al., 2017) in which the creation of an automated tool removes incentives for people to continue creating the training texts that power that LLMs. Encourage users to re-engage with the data source (e.g., by posting their own experiences to an online community).
5. Discussion
5.1. Perceptions of risks and benefits of LLMs
@@
Participants reported largely positive attitudes towards NLP tools with important caveats. For example, workshop participants expressed high hopes about the development and adoption of NLP tools in healthcare but unanimously agreed that NLP tools for maternal health should always be designed to be assistive rather than autonomous (holding on to the human). Settings in which NLP tools were viewed to be particularly useful to healthcare providers included: (i) reducing administrative burden, including better/faster communication and coordination of care across medical departments, and (ii) improving medical education and training, including assisting staff in developing better cultural sensitivity. On the other hand, tools for decision-making settings were viewed negatively. These results contrast with current trends in NLP research (§1), which tend towards surveillance of birthing people and tasks like risk prediction.
@@
Birthing people viewed NLP tools very positively, especially for providing information and recommendations (centering their agency and autonomy). For example, participants wrote that they wished they had the chatbot during their pregnancy: “I wish it [(AI)] had been around when my son was a newborn so I could interact with it during late night feedings. One, to give me something to do, and two, to make me feel like I wasn’t alone” (B12). In particular, birthing people emphasized how comforting it would be to have a fast and convenient resource to assuage fears about whether what they were experiencing was normal: “Just to have something there to ask questions to when I am not sure as to what is happening or when I need a quick answer” (B9). Some participants cautioned that such tools should provide additional context to help patients understand whether the suggestions or advice are applicable to their specific situation due to differing levels of medical understanding.
These positive impressions need to be weighed carefully against research demonstrating potential harms. On one hand, both birthing people and clinicians are desperately in need of support, as e.g., postpartum depression rates (Stewart and Vigod, 2016) and clinician burnout rates (Dzau et al., 2018) demonstrate. Challenges around healthcare costs, inaccessibility, and community distrust of the medical establishment (Cuevas et al., 2016) support the positive views of our participants towards NLP and LLM-based tools. But on the other hand, the current capabilities of these tools are limited and their vulnerabilities well-documented (Weidinger et al., 2021), and for some challenges, non-technical solutions already exist whose implementation might be further delayed by allocating resources to NLP technologies. Researchers and practitioners should be ready to abandon NLP technologies for maternal health if there are not clear benefits.
5.2. LLMs as part of a larger ecosystem to meet information needs